[ A student traveled 194 km in 20 days. at the same ratio, how far would the student travel in 80 days. ]
Answer by TimothyLamb(4379)

---
164/20 = d/t
d = (164/20)*t
d = (164/20)*80
---
answer:
d = 656 km
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 4-equations 4-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php